OpenStudy (anonymous):

find all solutions tan^2 θ − 1 = 0

OpenStudy (anonymous):

\(\large tan^2\theta-1=0 \) \(\large tan^2\theta=1 \) \(\large tan\theta=1 \) or \(\large tan\theta=-1 \) can you solve these two equations?

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

The person before me has made it simple. You find the value of theta for tan inverse 1 and tan inverse -1.

OpenStudy (anonymous):

how do you do that

OpenStudy (anonymous):

have you used the unit circle before?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok... let's just do that first equation, \(\large tan\theta=1 \) in terms of sine and cosine, that equation can be written as \(\large \frac{sin\theta}{cos\theta}=1 \). what angle on the unit circle gives the same value for the sin and cos ?

OpenStudy (anonymous):

pi/4 and 3pi/4

OpenStudy (anonymous):

pi/4 is one of 'em but 3pi/4 are opposites (it is a solution for the second equation though) see page 3 here: http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (anonymous):

5pi/4

OpenStudy (anonymous):

yes... so for the first equation, the solutions are pi/4, and 5pi/4 what about \(\large tan\theta=-1 \)... can you give me the solution(s) for this one?

OpenStudy (anonymous):

3pi/4 and7pi/4

OpenStudy (anonymous):

yes... those are the solutions for theta from 0 to 2pi. did you want the general solution?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

oops... ALL solutions... ok...

OpenStudy (anonymous):

sorry... the solutions are multiples of pi/2 apart

OpenStudy (anonymous):

Because ByteMe is doing his best to help, here's a medal.

OpenStudy (anonymous):

so your general solution (all solutions) is pi/4 + n*pi/2, where n = ....-2, -1, 0, 1, 2, ...

OpenStudy (anonymous):

thanks @CeroOscura94 ... :)

OpenStudy (anonymous):

thank i got it :)